CHILD'S PLAY COMPANY (A)

Child's Play Company makes a plastic rattle for toddlers. The rattle is generally marketed through exclusive retailers located in upscale shopping malls. The company used 23 retailers. In late 2016, Diana Suares, the president of the company, was considering alternative marketing planspresented to her by Bill Dings, the marketing manager. Based on sales from January through October 2016, Diana expected that 2016 sales would amount to 300,000 units at $8.00 per unit. Diana Suares also had with her some cost data for 2016 supplied by the CFO, Don Capp. Don expects that these costs are reliable estimates for a volume up to 400,000 units. Beyond 400,000 units, the company would have to rent additional machines (with a capacity of 100,000 units each)at an annual cost of $50,000 per machine (in 2016). The data are presented in Exhibit A. Total fixed manufacturing and selling and administrative costs are each expected to increase in 2017 by 10 percent because of inflation. Variable costs per unit and the selling price would stay the same as in 2016

Exhibit A (2016 Cost Data)

Manufacturing costs for rattles (based on production volume of 300,000 units).

Direct material $0.80 per unit *

Direct labor $10 per hour *

(Each worker can make 20 units in an hour)

Packaging $0.75 per unit *

Power, supplies, indirect labor, and other variable production costs $1.20 per unit *

Supervisory salaries, equipment rental, and miscellaneous production costs $1.80 per unit

* These costs vary in total with production volume

Selling and administrative costs (based on sales volume of 300,000 units)

Commissions to sales staff 10% of price to retailer **

Shipping $0.50 per unit **

Advertising and promotion $0.60 per unit

Administrative staff salaries, depreciation on office equipment, etc $0.90 per unit

** These costs vary in total with sales volume.

Required For 2017

1. What is the materials cost at 300,000 units?

2. What is the materials cost at 400,000 units?

3. What is the total variable cost at 200,000 units?

4. What is the total variable cost at 400,000 units?

5. What is the fixed manufacturing cost at 300,000 units?

6. What is the fixed manufacturing cost at 400,000 units?

7. What is the fixed manufacturing cost at 450,000 units?

8. What is the total fixed cost at 300,000 units?

9. What is the total fixed cost at 400,000 units?

10. What is the contribution margin per unit?

11. What is the contribution margin if 1000 units are sold?

12. What is the contribution margin if 100,000 units are sold?

13. What is the profit at the break even volume?

14. What is the contribution margin at the break even volume?

15. What is the break even volume?

16. What is the profit at 1000 units above break even?

17. What is the profit at 2000 units below the break even?

18. What is the profit at 100,000 units above break even?

19. What is the margin of safety at 1000 units above break even?

20. What is the sales volume required to give $100,000 profit before tax?

21. What is the sales volume required to give $180,000 after tax, assuming a tax rate of 40%
Answer #1

As a rule, only the first 4 sub-parts can be answered, unless other specific sub-parts have been requested (subjected to max 4). As no specific sub-parts are requested, the first 4 are solved.

1. What is the materials cost at 300,000 units?
Direct material cost per unit = $0.80
Total no. of units = 300,000
Therefore, total material cost = 0.80 * 300,000 = 240000

2. What is the materials cost at 400,000 units?
Direct material cost per unit = $0.80
Total no. of units = 400,000
Therefore, total material cost = 0.80 * 400,000 = 320000

3. What is the total variable cost at 200,000 units?
Let us list all the variable costs:

Particulars Amount ($) per unit
1 Direct Material 0.8
2 Direct Labour # 0.5
3 Packaging 0.75
4 Power, supplies, indirect labor, and other variable production costs 1.2
5 Commissions to sales staff 0.8
( 10% of price to retailer i.e $8)
6 Shipping 0.5
Therefore, total variable cost 4.55
# Direct Labour
Cost per hour $10
Number of units per hour 20
Therefore, cost per unit =
Cost per hour / Number of units per hour $0.5

As from above, total variable cost = 4.55
Total no. of units = 200,000
Therefore, total variable cost = 4.55 * 200,000 = $910000

4. What is the total variable cost at 400,000 units?

As from above, total variable cost = 4.55
Total no. of units = 400,000
Therefore, total variable cost = 4.55 * 400,000 = $1820000